# Index Terms: Machine-generated code detection, large language models, code generation, empirical analysis

### 24 Jan 2024 v2

@@ 5.2 Evaluation Metric
5.3 Baselines and Models
Our evaluation is benchmarked against a diverse range of zero-shot machine-generated text detection techniques, all of which leverage the predicted token-wise conditional distributions of the source model for detection. A supervised baseline is also included to demonstrate the effectiveness of our method.
@@ 5.3 Baselines and Models
As for models to generate machine’s code, in consideration of GPU memory limitations, we judiciously selected a range of advanced LLMs with varying parameter counts, from 1 billion to 7 billion. We selected Incoder [24], Phi-1 [50], StarCoder [51], WizardCoder [25], CodeGen2 [4] and CodeLlama [6] from Huggingface  [^3] https://huggingface.co/models. This diverse array of models, each trained on distinct datasets and boasting different capacities, offers a comprehensive spectrum for evaluating our method.
5.4 Experimental Setup
For code generation with different models, we adopted the top-$p$ sampling strategy [41] with $p$=$0.95$ following [1]. We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section 3.2. The maximum length constraint for generated codes was set at $128$ tokens. With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model [3], revealed optimal values. Consequently, we set $\alpha$ and $\beta$ to $0.5$, while $\lambda_{\text{spaces}}=3$ and $\lambda_{\text{newlines}}=2$. For all experiments, we maintained a consistent configuration of generating $50$ perturbations. For the DetectGPT and DetectLLM, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model [52]. And as for the supervised baseline GPTSniffer, we trained the CodeBERT model for 5 epochs with a batch size of 16 and a learning rate of 2e-5 with AdamW optimizer. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
@@ 5.5 Detection Performance (RQ1)
It is also worth mentioning that our method, leveraging a zero-shot framework, frequently surpasses the accuracy of the supervised GPTSniffer. The unstable performance of the supervised model reveals that it’s hard to detect machine-generated codes relying on the training data. The ability of our zero-shot approach to adapt and retain high performance under different levels of generative temperature variations is indicative of its sophisticated detection capability, making it a valuable tool for maintaining the integrity of codebases against the infiltration of machine-generated code in practical applications.
@@
To further illustrate the effectiveness of our approach, we present some representative examples of machine-generated code detection by DetectCodeGPT and two most competitive baselines in Figure 5. We set the mean score of all the code snippets as the decision threshold $\epsilon$ for each method.
Examples 1 is a machine-generated code snippet. DetectCodeGPT can correctly detect it, while the baselines fail. In this example, the first and last “if” statements are separated from the rest of the code with newlines, and the two “if” statements in the middle are modularized together since they have similar functionalities. Such modularization and separation of code blocks are captured by DetectCodeGPT thanks to the stylized perturbation. Example 2 and 3 are human-written codes that are misclassified by other baselines, but DetectCodeGPT correctly identifies them. In Example 2, we can observe that the code blocks are sometimes separated with newlines (e.g., lines 5-7), but sometimes not (as seen with the rest of the code). In Example 3, although the code blocks are well separated with newlines, the human author omitted the spaces between operators “//” and “/”, but add spaces between “*” and “+”. Such freely and randomly stylized code reveals the inherent randomness in human coding habits. It’s hard for the baselines that only rely on token-wise conditional distributions to capture such randomness in coding styles. In contrast, DetectCodeGPT effectively leverages the style information to distinguish between machine-generated and human-written code again. These examples demonstrate the effectiveness of our method in detecting machine-generated code via stylized code perturbation.
However, there are also some cases where DetectCodeGPT fails to detect machine-generated code. Example 4 is a human-written code misclassified as machine-generated by all the approaches. We can observe that this code snippet is well-structured with newlines and spaces, and the comments are well-aligned with the code. Its resemblance to machine-generated code is striking, posing a significant challenge for distinction. This example highlights the difficulty of detecting machine-generated code among well-structured human-written codes with standard coding styles.
5.6 Ablation Study (RQ2)
@@
In our ablation study, we compared the effectiveness of different perturbation strategies for detecting machine-generated code using the CodeLlama (7B) model on The Stack dataset. The results summarized in Table V illuminates the comparative advantage of our stylistic perturbation approach. We observe that both newline and space perturbations independently offer substantial improvements over the traditional MLM-based (CodeT5+) perturbation technique as in DetectGPT and DetectLLM [11, 45] for natural language. Also, the combination of newline and space perturbations further enhances the detection performance, with the highest AUROC score of 0.8852 at $T=0.2$ and 0.6660 at $T=1.0$. The consistent outperformance of our combined perturbation strategy across both temperature settings affirms its potential as a robust solution for detecting machine-authored code.
5.7 Impact of Perturbation Count (RQ3)
To gauge the impact of perturbation count on the efficacy of our method, we conducted experiments with varying numbers of perturbations.
@@ 5.7 Impact of Perturbation Count (RQ3)
Results in Table VI reveal a rapid ascent in the AUROC score as the number of perturbations increases from 10 to 20, underscoring the efficiency of our perturbation approach. Notably, an increase to 20 perturbations already yields robust detection performance, with further increments leading to diminishing improvements. This suggests that our method requires a relatively small number of perturbations to effectively discern between human- and machine-authored code. This implies that our method is not only effective but also resource-efficient.
5.8 Performance of Cross-Model Detection (RQ4)
@@
In previous sections, we primarily assessed the efficacy of DetectCodeGPT within a white-box framework, where the detection model can be the same as the source model. Nonetheless, in real-world applications, it is often infeasible to access to the original code generation model for detection. To simulate such conditions, we engaged in cross-model detection experiments, wherein distinct models were utilized to compute the naturalness scores for detecting code generated by others. These evaluations were carried out using The Stack dataset at a temperature of 0.2.
The results presented in Figure 6 highlight DetectCodeGPT’s adaptability in cross-model detection. While the algorithm excels in the white-box setting, its performance endures with only a slight reduction when subjected to cross-model application. For instance, StarCoder, when detecting code generated by WizardCoder and CodeLlama, yields AUROC scores of 0.92 and 0.87, respectively, compared to an AUROC of 0.93 when detecting its own output. And we can also notice a performance decrease when detecting CodeGen2’s output. This is possibly due to the fact that CodeGen2 is trained on a more diverse dataset containing more natural language text [4]. However, Phi-1 demonstrates a relative proficiency with a score of 0.83 to detect CodeGen2’s output, which implies an ensemble of diverse detection models may enhance the system’s robustness, as suggested in [53].
@@
6 Discussion
6.1 Why is DetectCodeGPT Effective?
@@
6.1.1 Preservation of Code Correctness
While the perturbation method in DetectGPT masks continuous spans of text and requires another LLM to recover, such recovery often leads to structurally compromised code. The unique syntax and constraints inherent to programming languages mean that even minor misinterpretations (e.g., misuse of an identifier) can render the code non-functional. The originally masked spans could be the only correct solution to recover the code and the recovered version from LLM bring minor mistakes easily. Such code-cracking perturbations will possibly lead to both a decrease in naturalness scores for human’s and machine’s codes, violating the assumption that the perturbation will have minimal impact on the code’s naturalness score if it is human-written in Section 2.2. In contrast, inserting newlines and spaces does not affect the correctness of the code in most cases. So our perturbation strategy ensures that the underlying structure and correctness of the code remain intact, thereby ensuring the effectiveness of our method.
6.1.2 Emulation of Human Randomness
@@
As discussed in Section 3.4.5, human inherently exhibit less naturalness and more randomness in their use of stylistic tokens such as spaces and newlines than machines. For example, a human programmer may freely insert whitespace, especially newlines, in the code as they deem fit, whereas a machine programmer usually tries to stylize the code in a more standardized and modularized manner. Our proposed perturbation strategy mimics human’s free usage of spaces and newlines, thereby making the perturbation more “random” as is desired according to Section 3.
6.2 Strength of DetectCodeGPT
DetectCodeGPT presents a perturbation strategy that is not only efficient but also maintains the structural integrity of the code while closely mirroring human coding styles. We summarize the strengths of DetectCodeGPT as follows:
6.2.1 Enhanced Efficiency
Since it’s necessary to perturb each text for multiple times to ensure the effectiveness of detection [11, 45], as verified in Section 5.7, perturbing code using masked language modeling with other pretrained models will lead to huge computational cost. On the contrary, our method will save much more time and bring more efficiency. This translates to a significant conservation of computational resources, thus expediting the detection process and enhancing scalability. Mathematically, if $T_{\text{DetectGPT}}$ and $T_{\text{Ours}}$ denote the computational times of DetectGPT and our method respectively, we know that $T_{\text{Ours}}\ll T_{\text{DetectGPT}}$.
6.2.2 Superior Zero-Shot Performance
@@
DetectCodeGPT distinguishes itself with a zero-shot learning capability, enabling it to detect machine-generated code without the necessity for training on extensive datasets. This model-agnostic advantage means that it can be generalized across various code LLMs and can still remain robust even when applied to unseen models, as demonstrated and verified in Section 5.8. In comparison to the supervised methods, which may suffer from data drift or the need for constant updates, our approach is not only more adaptable but also delivers better performance across various data sources. This makes DetectCodeGPT an agile and reliable tool for machine-generated code detection in a variety of contexts.
6.3 Limitations and Future Directions
@@
The main limitations of our work lie in the following two aspects: Firstly, due to the time and computational constraints, we only focus on a set of LLMs within 7B parameters for both empirical analysis and the machine-generated code detection task. As the landscape of LLMs rapidly evolves, incorporating a wider array of more and larger LLMs could significantly bolster the generalizability and robustness of our findings.
@@
Secondly, our current analysis centers exclusively on Python code, while the features of other programming languages may not be fully explored and the effectiveness of our method on other languages is not verified. However, based on the analysis of our method in Section 6.2, we believe that our method can be effectively generalized to other programming languages, especially where the functionality of code won’t be much affected after inserting newlines and spaces like C/C++, Java, and JavaScript.
@@
Looking ahead to future work, we plan to further improve the effectiveness of DetectCodeGPT when detecting machine-generated code at higher levels of generation randomness, such as $T$=$1.0$. We note from Table IV that although DetectCodeGPT outperforms other baselines at $T$=$1.0$, there is still large room for improvement. Although ensembling multiple detection models may help to improve the detection performance [53], we look forward to exploring more effective perturbation strategies based on code style to further enhance the detection efficacy.
Additionally, the development of watermarking strategies [54] for machine-generated code represents an intriguing avenue for future research. Given that our method can effectively identify machine’s code, it could be adapted to assist in the creation of watermarking techniques. For instance, embedding a watermark could be achieved by inserting newlines and spaces in a code-specific pattern, which would increase detectablility without compromising functionality.
7 Related Work
@@
However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-produced text [13]. GPTSniffer was first proposed to detect machine-generated code with supervised CodeBERT training [47]. And the predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training of the model or at the time of code generation [13, 54]. The detection of these watermarks subsequently enables the recognition of code generated by machines. It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism [62, 63]. Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.
8 Conclusion
In this paper, we have explored the nuanced differences between machine-generated and human-written code to enhance the trustworthiness of software development in the era of AI. By pioneering a comparative analysis across three aspects of code, namely, naturalness, style, and correctness, we have provided new insights into code provenance that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylized tokens such as whitespaces. Our proposed detection method, DetectCodeGPT, capitalizes on these insights, specifically targeting the stylized nuances of code with a novel perturbation strategy that is simple yet effective. The empirical success of DetectCodeGPT against several benchmarks confirms its potential to bolster security measures in programming and establish a more symbiotic relationship between human developers and LLMs. Our work marks a significant step towards responsible AI use in software engineering, ensuring clarity in code authorship and maintaining the integrity of code as a craft.
Data Availability
All the experimental data and source code used in this work are available at https://github.com/YerbaPage/DetectCodeGPT.

### 24 Mar 2024 v3

@@ V-C Evaluation Metric
V-D Baselines
Our evaluation is benchmarked against a diverse range of zero-shot machine-generated text detection techniques. And a supervised baseline is also included to demonstrate the advantages of our zero-shot method:
@@ V-D Baselines
V-E Experimental Setup
For code generation with different models, we adopted the top-$p$ sampling strategy [36] with $p$=$0.95$ following [1]. We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section III-B. The maximum length constraint for generated code was set at $128$ tokens. With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model [3], revealed optimal values. Consequently, we set $\alpha$ and $\beta$ to $0.5$, while $\lambda_{\text{spaces}}$=3 and $\lambda_{\text{newlines}}$=2. For all experiments, we maintained a consistent configuration of generating $50$ perturbations. For the DetectGPT and DetectLLM, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model [48]. And as for the supervised baseline GPTSniffer, we trained the CodeBERT model for 5 epochs with a batch size of 16 and a learning rate of 2e-5 with AdamW optimizer. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
@@ V-F Detection Performance (RQ1)
It is worth mentioning that our zero-shot framework often outperforms the supervised GPTSniffer, highlighting the challenges of detecting machine-generated code with training data-dependent supervised models. The stable performance across various generation settings showcases our method’s advanced detection capabilities. This makes it an effective solution in practical applications.
To further illustrate the effectiveness of our approach, we present some representative examples by DetectCodeGPT and two most competitive baselines in Figure 5, using decision thresholds $\epsilon$ set at the mean scores of all code snippets.
Examples 1 is a machine-generated code snippet. DetectCodeGPT can correctly detect it, while the baselines fail. In this example, the first and last “if” statements are separated from the rest of the code with newlines, and the two “if” statements in the middle are modularized together since they have similar functionalities. Such modularization and separation of code blocks are captured by DetectCodeGPT thanks to the stylized perturbation. Example 2 and 3 are human-written code that are misclassified by other baselines, but DetectCodeGPT correctly identifies them. In Example 2, we can observe that the code blocks are sometimes separated with newlines (e.g., lines 5-7), but sometimes not (as seen with the rest of the code). In Example 3, although the code blocks are well separated with newlines, the human author omitted the spaces between operators “//” and “/”, but add spaces between “*” and “+”. Such freely and randomly stylized code reveals the inherent randomness in human coding habits. Baselines, relying solely on token-wise conditional distributions, struggle to capture the coding style’s randomness, whereas DetectCodeGPT effectively utilizes style information to discern between machine-generated and human-written code, demonstrating its prowess in detection through stylized code perturbation.
However, there are also cases where DetectCodeGPT fails. Example 4 is a human-written code misclassified as machine-generated by all the approaches. We can observe that this code snippet is well-structured with newlines and spaces. Its resemblance to machine-generated code is striking, posing a significant challenge for distinction. This example highlights the difficulty of detecting machine-generated code among well-structured human-written code with standard coding styles.
@@
V-G Ablation Study (RQ2)
@@
In our ablation study, we compare the effectiveness of different perturbation strategies for detecting machine-generated code using the CodeLlama (7B) model on The Stack dataset. The results summarized in Table V illuminates the comparative advantage of our stylistic perturbation approach. We observe that both newline and space perturbations independently offer substantial improvements over the traditional MLM-based (CodeT5+) perturbation technique as in DetectGPT and DetectLLM [9, 41] for natural language. Also, the combination of newline and space perturbations further enhances the detection performance, with the highest AUROC score of 0.8852 at $T=0.2$ and 0.6660 at $T=1.0$. The consistent outperformance of our combined perturbation strategy across both temperature settings affirms its potential as a robust solution for detecting machine-authored code.
V-H Impact of Perturbation Count (RQ3)
To gauge the impact of perturbation count on the efficacy of our method, we conduct experiments with varying numbers of perturbations.
@@ V-H Impact of Perturbation Count (RQ3)
Results in Table VI reveal a rapid ascent in the AUROC score as the number of perturbations increases from 10 to 20, underscoring the efficiency of our perturbation approach. Notably, an increase to 20 perturbations already yields robust detection performance, with further increments leading to diminishing improvements. This suggests that our method requires a relatively small number of perturbations to effectively discern between human- and machine-authored code. This implies that our method is not only effective but also efficient.
V-I Performance of Cross-Model Code Detection (RQ4)
@@
In previous sections, we primarily assess the efficacy of DetectCodeGPT within a white-box framework, where we can get access to the logits of the original code generation model, as defined in Section IV-A. In real-world scenarios, accessing the original model for code generation detection is often impractical, so we conduct cross-model detection experiments, where we use other LLMs as surrogate models to compute the naturalness scores.
The evaluation results on The Stack dataset at $T=0.2$ presented in Figure 6 highlight DetectCodeGPT’s adaptability in cross-model detection. While the algorithm excels in the white-box setting, its performance endures with only a slight reduction in cross-model application. For instance, StarCoder, when detecting code generated by WizardCoder and CodeLlama, yields AUROC scores of 0.92 and 0.87, respectively, compared to an AUROC of 0.93 when detecting its own output. We can also notice a performance decrease when detecting CodeGen2’s output. This is possibly due to the fact that CodeGen2 is trained on a more diverse dataset containing more natural language text [4]. However, Phi-1 demonstrates a relative proficiency with a score of 0.83 to detect CodeGen2’s output, which implies an ensemble of diverse detection models may enhance the system’s robustness, as suggested in [49].
@@
VI Discussion
VI-A Why is DetectCodeGPT Effective?
@@
VI-A1 Preservation of Code Correctness
The mask-and-recover perturbation in DetectGPT brings minor mistakes easily (e.g., misuse of an identifier), rendering the code non-functional and has negative impact on the naturalness score. Such code-cracking perturbations will violate the assumption of minimal impact on the code’s naturalness score if it is human-written in Section II-B. In contrast, inserting newlines and spaces does not affect the correctness of the code in most cases, thereby ensuring the effectiveness of our method.
VI-A2 Emulation of Human Randomness
@@
As discussed in Section III-D5, human inherently exhibit less naturalness and more randomness in their use of stylistic tokens such as spaces and newlines than machines. For example, a human programmer may freely insert whitespace, especially newlines, in the code as they deem fit, whereas a machine programmer usually tries to stylize the code in a more standardized and modularized manner. Our proposed perturbation strategy mimics human’s free usage of spaces and newlines, thereby making the perturbation more “random” as is desired according to Section III.
VI-B Strength of DetectCodeGPT
@@
Compared with existing methods, DetectCodeGPT eliminates the need to perturb code multiple times for each LLM and thus brings more efficiency. Compared with supervised counterparts, DetectCodeGPT distinguishes itself with a zero-shot learning capability, enabling it to detect machine-generated code without the necessity for training on extensive datasets. This model-agnostic advantage means that it can be generalized across various code LLMs.
VI-C Limitations and Future Directions
@@
The main limitations of our work lie in the following two aspects: Firstly, due to the computational constraints, we only focus on a set of LLMs within 7B parameters. As the landscape of LLMs rapidly evolves, incorporating a wider array of more and larger LLMs could significantly bolster the generalizability and robustness of our findings.
@@
Secondly, our current analysis centers exclusively on Python code, while the features of other programming languages may not be fully explored. However, based on the analysis of our method in Section VI-B, we believe that our method can be effectively generalized to other languages, especially where the functionality of code won’t be much affected after inserting newlines and spaces like C/C++, Java, and JavaScript.
@@
Looking ahead to future work, we plan to further improve the effectiveness of DetectCodeGPT when detecting machine-generated code at higher levels of generation randomness. We note from Table IV that although DetectCodeGPT outperforms other baselines at $T$=$1.0$, there is still large room for improvement. Although ensembling multiple detection models may help improve the detection performance [49], we look forward to exploring more effective perturbation strategies based on code styles to further enhance the detection efficacy.
VII Related Work
@@
However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-generated text [10]. GPTSniffer was first proposed to detect machine-generated code with supervised CodeBERT training [45]. The predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training or generation [10, 58, 59]. The detection of these watermarks subsequently enables the recognition of code generated by machines. It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism [60, 61]. Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.
VIII Conclusion
In this paper, we perform an in-depth analysis of the nuanced differences between machine- and human-authored code across three aspects of code including lexical diversity, conciseness, and naturalness. The results provide new insights that machines tend to write more concise and natural code, adhering to common programming paradigms, and the disparity is more pronounced in stylized tokens such as whitespaces that represent the syntactic segmentation of code. Based on these insights, we have proposed a new detection method, DetectCodeGPT, which introduces a novel stylized perturbation strategy that is simple yet effective. The experimental results of DetectCodeGPT confirm its effectiveness, demonstrating its potential to help maintain the authorship and integrity of code.